OpenStudy (imnotgoodatmath):

Which is a power of 10?

OpenStudy (anonymous):

It's basic math , think about it ;)

OpenStudy (imnotgoodatmath):

10X10

OpenStudy (imnotgoodatmath):

?

OpenStudy (imnotgoodatmath):

0.001001 10 1,001 105

OpenStudy (vishweshshrimali5):

Try expressing each number as \(\large{10^{\text{whole number}}}\)

OpenStudy (vishweshshrimali5):

For example: 105 can `not` be expressed as 10^(whole number) Thus `105 is NOT the correct answer`

OpenStudy (vishweshshrimali5):

Can you express 1001 as 10^(whole number) ? HINT: 1000 = 10^3

OpenStudy (imnotgoodatmath):

sorry, its 10^5 and 1^0

OpenStudy (vishweshshrimali5):

no no no 1001 is `not` 10^5 and 1^0 In fact: `1001 can NOT be represented as 10^(whole number)`

OpenStudy (imnotgoodatmath):

so its the first one

OpenStudy (anonymous):

It's so easy bro!

OpenStudy (anonymous):

:P

OpenStudy (vishweshshrimali5):

No the second one

OpenStudy (vishweshshrimali5):

\[{10 = 10^1}\]